Many roofs are designed in the shape of a pyramid. The formula:
Area= 2s * the square root of 0.25s(squared) + h(sqaured) where s = the lenghth of a side of the sqaure base and h = the height of the base of the pyramid to the vertex.
solve the equation for h algebraically
suppose that you are tiling a roof that is a pyramid with a square base having sides of length 42.0 ft and the tile is sold in bundles that cover an area of 25 square feet. Determine algebraically the maximum height you could build the roof if you have 80 bundles of tile.

Answer by khwang(438)

As I understand, the area is
A(s,h) = 2s sqrt(0.25 s^2 + h^2)
When s =42,
A(h) = 84 sqrt(0.25 *42^2 + h^2) = 84 sqrt(441 + h^2)
So, h = sqrt((Area/84)^2 - 441)
= sqrt(Bundles/4410000 - 441), where Bundles = Area/25 ..integer
Since 80 bundles of tiles can cover 80*25 = 2000 ft^2.
We have h = sqrt((Area/84)^2 - 441) = 11.22 ft ... max height
Kenny